OpenStudy (anonymous):

Can someone explain to me, in complete sentences, the relationship between the equations of perpendicular lines?

OpenStudy (anonymous):

their slope are opposite reciprocal

OpenStudy (mathteacher1729):

The following lines are perpendicular because their slopes are negative reciprocals. (one is m, the other is -1/m. :) \[(y-y_0) = m(x-x_0)\] \[(y-y_0)= -\frac{1}{m}(x-x_0) \]